OpenStudy (anonymous):

A car travels at 35 mph. How long does it take for it to travel 96.25 mi.?

OpenStudy (raffle_snaffle):

v=d/t

OpenStudy (raffle_snaffle):

solve for t

OpenStudy (raffle_snaffle):

do you understand>

OpenStudy (anonymous):

no

OpenStudy (raffle_snaffle):

velocity=distance/total time

OpenStudy (raffle_snaffle):

we know d and v so solve for t

OpenStudy (anonymous):

d=35mph ?

OpenStudy (anonymous):

distance = speed x time but we are looking for time, so we can change the formula to : D/S = T D = 96.25 S = 35 now sub 96.25/35 = T 2.75 = T It would take 2.75 hrs or 2 3/4 hrs Does that make sense to you ?

OpenStudy (anonymous):

Oh... Thanks a lot! i was thinking D=35 and S= 96.25